CRY+VIS Interface Meeting
Date: 2021/9/24 13:00-
Participants: Takahashi, Uchiyama, Sato, Aso, Akutsu, Ushiba, Tomaru, Yasui, Yokozawa, Ikeda, Kimura
- IX: Health check. Investigating the disturbance. Effect due to the convection will be checked. Diagonalization of IP is not good.
- EX: Health check. Installation of ACCs will be scheduled after the health check.
- IY: H4 geophones for IY/EY were delivered. One of them will be used in OMC chamber temporally.
- EY: F0 FR should be modified.
- TM height setting: Summary of adjustable height range.
- Commissioning in air: Health check.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers:
- Heater control: Box covers for the adapter were installed.
Date: 2021/9/10 14:00-
Participants: Takahashi, Sato、Fabian, Uchiyama, Aso, Sugiyama, Terrence, Hirata
- Health check in IX and EX.
- We reported before a problem with a RM coil.
- Last week we assessed the problem was one of the three cable segments used.
- This week we we implemented a work around: instead of using the usual pins 2 and 7, we changed the pins to places 1 and 6 for segments 2 and 3.
- Although it's possible to switch back to pins 2 and 7 in the upper connector of sengment 2, the connector is in an ackward place and we preferred to leave it alone.
- Yasui-san and Ikeda-san provided an in-air robust adapter.
- We reported before that an OSEM coil had a slightly larger resistance than expected.
- This week we identified the source of the problem: the pins in a male connector are not making good contact.
- The connector is at the base of the preisolator and it's easy to access and fix.
- We'll fix it next week.
- There was an earthquake on September 19th that changed the pitch of the mirror.
- We aligned the mirror such that the oplev beam was again at the centre of the tilt QPD.
- However, V1 OSEM went out of range with 300 counts.
- A measurement of IM-Y transfer function suggests that the OSEM flags are not touching their bodies.
- Before adjusting the position of the V1 OSEM, we decided to try to bring it into range by tilting the BF and IM and adjusting the height of the BF.
- We replaced a broken flip adapter keeping one of the BF picomotors from working.
- Type-A team: Health check in IX and EX. Assembly of H4 geophones for IY and EY.
- Type-B team: BF picomotor in PR3, OSEM coil in SR3, BS outerframe.
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
Effectiveness of ”BS outer frame" reinforcement. rev_BS outer frame(recalculation).pdf
The week from 9/27
- Takahashi: 9/27-10/1
- Sato: 9/27-10/1
- Hirata: 9/27-10/1
- Yano: 9/27-10/1
The week from 10/4
- Takahasdhi: 10/4-8
- Sato: 10/4-8
- Hirata: 10/4-8
- Yano: 10/4-8